# Benchmarking Retrieval-Augmented Generation for Chemistry

### v1

@@ 1 Introduction
Retrieval-augmented generation (RAG) (Gao et al., 2023) has emerged as a powerful paradigm for enhancing large language models (LLMs) with external knowledge sources. By incorporating retrieval into the generation process, RAG can effectively mitigate hallucinations (Zhang et al., 2023b) and inject up-to-date domain-specific information into LLMs (Siriwardhana et al., 2023). These capabilities are particularly valuable in scientific domains, where factual accuracy and timely knowledge are critical. A typical scientific RAG system consists of two components: (1) a retriever that selects relevant documents or facts from a scientific knowledge base, and (2) a generator, often an LLM, that integrates the retrieved content to produce informed and coherent responses. Such frameworks have shown promising applications in domains like biomedicine (Xiong et al., 2024).
In chemistry, LLMs have shown remarkable potential across various tasks, including molecular captioning (Li et al., 2024), chemical reasoning (Tang et al., 2025), and reaction prediction (Shi et al., 2023). However, chemistry is a highly specialized and dynamic discipline, characterized by complex terminologies, domain-specific conventions, and rapidly-evolving knowledge. As a result, LLMs trained on general corpora often fail to generate grounded and accurate responses, instead producing hallucinated or outdated content (Zhang et al., 2023a; Wang et al., 2024b). RAG presents a natural solution to these limitations, allowing models to retrieve and incorporate trusted chemical knowledge during inference.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) enhances large language models by incorporating external knowledge sources (Lewis et al., 2020). It has been shown to reduce hallucinations (Ayala & Bechard, 2024) and provide access to up-to-date information (Fan et al., 2024). Recent work has sought to improve RAG performance through various enhancements, including more effective retrieval mechanisms (Glass et al., 2022), iterative retrieval-and-reasoning pipelines (Trivedi et al., 2022; Jin et al., 2025), and the integration of long-context language models to better handle extended inputs (Jin et al., 2024). While substantial progress has been made in general-domain RAG benchmarks (Asai et al., 2023; Yu et al., 2024; Kwiatkowski et al., 2019; Yang et al., 2018), relatively little attention has been given to the scientific domain. Although recent efforts, such as Xiong et al. (2024), begin to explore this direction in the medicine domain, the application of RAG to the chemistry domain remains underdeveloped. Notably, TextReact (Qian et al., 2023) applies text retrieval to tasks like reaction condition recommendation and one-step retrosynthesis. ChemLit-QA (Wellawatte et al., 2024) introduces a dataset for chemistry-oriented RAG, but its questions are generated from isolated paper excerpts and may lack real-world utility. Importantly, there remains a gap in the availability of high-quality, domain-specific corpora and comprehensive RAG benchmarks tailored to chemistry.
@@ 4 The ChemRAG-Toolkit
In ChemRAG-Toolkit, we select four representative retrievers for the retrieval process in RAG: BM25 (Robertson & Zaragoza, 2009), Contriever (Izacard et al., 2022), SPECTER (Cohan et al., 2020), and e5 (Wang et al., 2024a). In addition, we implement Reciprocal Rank Fusion (RRF, Cormack et al. (2009)) to combine the results from different retrievers.
@@ 5.1 Comparison of Backbone LLMs
PubChem BM25 47.19 12.98 36.00 27.73 30.98
Contriever 48.18 10.02 39.50 29.72 31.86
SPECTER 49.83 9.98 36.75 26.80 30.84
e5 46.86 8.61 40.50 30.65 31.66
RRF 48.84 9.08 37.38 29.58 31.22
PubMed BM25 46.86 12.02 38.63 28.14 31.41
Contriever 49.17 10.37 37.13 27.51 31.05
SPECTER 47.19 9.08 36.63 27.69 30.15
e5 46.53 10.36 39.63 25.07 30.40
RRF 48.18 8.98 37.13 25.70 30.00
USPTO BM25 49.50 11.55 44.00 56.17 40.31
Contriever 49.50 10.92 42.25 37.40 35.02
SPECTER 47.85 9.44 37.00 31.71 29.00
e5 47.52 11.05 38.13 37.55 33.56
RRF 49.17 10.70 43.38 56.68 39.98
Semantic Scholar BM25 45.54 7.18 37.25 29.72 29.92
Contriever 47.85 12.65 38.88 31.73 32.78
SPECTER 49.17 10.45 37.00 26.44 30.77
e5 45.21 10.78 38.75 31.52 31.57
RRF 44.55 8.91 39.13 31.76 31.09
OpenStax BM25 50.17 10.04 37.88 28.34 31.61
Contriever 49.50 12.66 36.38 27.95 31.62
SPECTER 50.50 11.88 37.13 28.20 31.93
e5 49.83 11.57 38.5 29.96 32.47
RRF 52.48 11.55 37.25 29.35 32.66
Wiki BM25 49.17 8.06 38.75 27.67 30.91
Contriever 48.84 10.33 37.25 29.14 31.39
SPECTER 47.52 9.21 39.25 27.22 30.8
e5 50.83 8.93 37.13 29.54 31.61
RRF 50.17 10.70 38.00 27.66 31.63
Chem -RAG Corpus BM25 49.83 6.51 38.13 34.99 32.37
Contriever 53.46 12.58 42.63 42.08 37.69
SPECTER 48.18 8.57 41.63 32.35 32.69
e5 47.19 7.56 37.13 42.24 33.53
RRF 52.81 8.80 44.5 45.83 37.99
@@ 8 Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329, and Apple PhD Fellowship. This research was also supported in part by the Division of Intramural Research (DIR), National Library of Medicine (NLM), National Institutes of Health (NIH). Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government.
@@ Appendix A Evaluation Metrics for Molecules
For molecular generation, we begin by verifying the validity of generated molecules using RDKit (Landrum et al., 2013) and then compute their exact match with reference solutions. However, a single textual description can correspond to multiple molecular structures, making exact matching a limited evaluation criterion. Moreover, expecting an LLM—even one fine-tuned with LoRA on specific instructions—to consistently generate outputs that perfectly match reference molecules is often unrealistic.
To address these challenges and provide a more comprehensive evaluation, we incorporate molecular similarity metrics, including similarity scores based on RDKit, MACCS, and Morgan fingerprints (Tanimoto, 1958; Schneider et al., 2015; Durant et al., 2002), alongside Levenshtein (Li & Liu, 2007) and BLEU scores.

### v2

@@ 1 Introduction
Retrieval-augmented generation (RAG) (Gao et al., 2023) has emerged as a powerful paradigm for enhancing large language models (LLMs) with external knowledge sources. By incorporating retrieval into the generation process, RAG can effectively mitigate hallucinations (Zhang et al., 2023a) and inject up-to-date domain-specific information into LLMs (Siriwardhana et al., 2023). These capabilities are particularly valuable in scientific domains, where factual accuracy and timely knowledge are critical. A typical scientific RAG system consists of two components: (1) a retriever that selects relevant documents or facts from a scientific knowledge base, and (2) a generator, often an LLM, that integrates the retrieved content to produce informed and coherent responses. Such frameworks have shown promising applications in domains like biomedicine (Xiong et al., 2024).
In chemistry, LLMs have shown remarkable potential across various tasks, including molecular captioning (Li et al., 2024), chemical reasoning (Tang et al., 2025), and reaction prediction (Shi et al., 2023). However, chemistry is a highly specialized and dynamic discipline, characterized by complex terminologies, domain-specific conventions, and rapidly-evolving knowledge. As a result, LLMs trained on general corpora often fail to generate grounded and accurate responses, instead producing hallucinated or outdated content (Zhang et al., 2023b; Wang et al., 2024b). RAG presents a natural solution to these limitations, allowing models to retrieve and incorporate trusted chemical knowledge during inference.
@@ 2.1 Retrieval-Augmented Generation
Retrieval-augmented generation (RAG) enhances large language models by incorporating external knowledge sources (Lewis et al., 2020). It has been shown to reduce hallucinations (Ayala and Bechard, 2024) and provide access to up-to-date information (Fan et al., 2024). Recent work has sought to improve RAG performance through various enhancements, including more effective retrieval mechanisms (Glass et al., 2022), iterative retrieval-and-reasoning pipelines (Trivedi et al., 2022; Jin et al., 2025), and the integration of long-context language models to better handle extended inputs (Jin et al., 2024). While substantial progress has been made in general-domain RAG benchmarks (Asai et al., 2023; Yu et al., 2024; Kwiatkowski et al., 2019; Yang et al., 2018), relatively little attention has been given to the scientific domain. Although recent efforts, such as Xiong et al. (2024), begin to explore this direction in the medicine domain, the application of RAG to the chemistry domain remains underdeveloped. Notably, TextReact (Qian et al., 2023) applies text retrieval to tasks like reaction condition recommendation and one-step retrosynthesis. ChemLit-QA (Wellawatte et al., 2024) introduces a dataset for chemistry-oriented RAG, but its questions are generated from isolated paper excerpts and may lack real-world utility. Importantly, there remains a gap in the availability of high-quality, domain-specific corpora and comprehensive RAG benchmarks tailored to chemistry.
@@ 4 The ChemRAG-Toolkit
In ChemRAG-Toolkit, we select four representative retrievers for the retrieval process in RAG: BM25 (Robertson and Zaragoza, 2009), Contriever (Izacard et al., 2022), SPECTER (Cohan et al., 2020), and e5 (Wang et al., 2024a). In addition, we implement Reciprocal Rank Fusion (RRF, Cormack et al. (2009)) to combine the results from different retrievers.
@@ 5.1 Comparison of Backbone LLMs
[…]
@@ 8 Acknowledgments
Research was supported in part by US DARPA INCAS Program No. HR0011-21-C0165 and BRIES Program No. HR0011-24-3-0325, National Science Foundation IIS-19-56151, the Molecule Maker Lab Institute: An AI Research Institutes program supported by NSF under Award No. 2019897, the Institute for Geospatial Understanding through an Integrative Discovery Environment (I-GUIDE) by NSF under Award No. 2118329, and Apple PhD Fellowship. This research was also supported in part by the Division of Intramural Research (DIR), National Library of Medicine (NLM), National Institutes of Health (NIH). Any opinions, findings, and conclusions or recommendations expressed herein are those of the authors and do not necessarily represent the views, either expressed or implied, of DARPA or the U.S. Government. This research has used the Delta/DeltaAI advanced computing and data resource, a joint effort of the University of Illinois Urbana-Champaign and its National Center for Supercomputing Applications, supported by the National Science Foundation (award OAC 2320345) and the State of Illinois. Access to DeltaAI was provided in part by the Illinois Computes project which is supported by the University of Illinois Urbana-Champaign and through allocation #250851 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which is supported by National Science Foundation grants #2138259, #2138286, #2138307, #2137603, and #2138296. The views and conclusions contained in this paper are those of the authors and should not be interpreted as representing any funding agencies.
@@ Appendix A Evaluation Metrics for Molecules
For molecular generation, we begin by verifying the validity of generated molecules using RDKit (Landrum and others, 2013) and then compute their exact match with reference solutions. However, a single textual description can correspond to multiple molecular structures, making exact matching a limited evaluation criterion. Moreover, expecting an LLM—even one fine-tuned with LoRA on specific instructions—to consistently generate outputs that perfectly match reference molecules is often unrealistic.
To address these challenges and provide a more comprehensive evaluation, we incorporate molecular similarity metrics, including similarity scores based on RDKit, MACCS, and Morgan fingerprints (Tanimoto, 1958; Schneider et al., 2015; Durant et al., 2002), alongside Levenshtein (Li and Liu, 2007) and BLEU scores.